please slove for the specified variable.
a= 3b-c for b

Answer by THANApHD(104)

a= 3b-c
3b-c=a
+c=> 3b-c+c = a+c
3b = a+c

*1/3=> 1/3 * 3b = (a+c) *1/3
b = (a+c)/3